# Learning multi-modal generative models with permutation-invariant encoders and tighter variational bounds

### v2

@@ 2 A tighter variational bound with arbitrary modality masking
This is simply a standard variational lower bound (Jordan et al., 1999; Blei et al., 2017) restricted to the subset $\mathcal{S}$ for $\beta=1$, and therefore $\mathcal{L}_{\mathcal{S}}(x_{\mathcal{S}},\theta,\phi,1)\leq\log p_{\theta}(x_{\mathcal{S}})$. To obtain a lower bound on the log-likelihood of all modalities, we introduce an (approximate) conditional lower bound
@@
For some fixed density $\rho$ on $\mathcal{P}(\mathcal{M})$, we suggest the overall bound
@@
which is a generalization of the bound suggested in Wu and Goodman (2019) to an arbitrary number of modalities. This bound can be optimized using standard Monte Carlo techniques, for example, by computing unbiased pathwise gradients (Kingma and Ba, 2014; Rezende et al., 2014; Titsias and Lázaro-Gredilla, 2014) using the reparameterization trick. For variational families such as Gaussian mixtures [^2] For MoE aggregation schemes, Shi et al. (2019) considered a stratified ELBO estimator as well as a tighter bound based on importance sampling, see also Morningstar et al. (2021), that we do not pursue here for consistency with other aggregation schemes that can likewise be optimized based on importance sampling ideas., one can employ implicit reparameterization (Figurnov et al., 2018). It is straightforward to adapt variance reduction techniques such as ignoring the scoring term of the multi-modal encoding densities for pathwise gradients (Roeder et al., 2017), see Algorithm 1 in Appendix H for pseudo-code. Nevertheless, a scalable approach requires an encoding technique that allows to condition on any masked modalities with a computational complexity that does not increase exponentially in $M$.
@@ 2.1 Multi-modal distribution matching
Likelihood-based learning approaches aim to match the model distribution $p_{\theta}(x)$ to the true data distribution $p_{d}(x)$. Variational approaches achieve this by matching in the latent space the encoding distribution to the true posterior as well as maximizing a tight lower bound on $\log p_{\theta}(x)$, see Rosca et al. (2018). We show similar results for the multi-modal variational bound. Consider, therefore the densities
@@
for the aggregated prior (Makhzani et al., 2016; Hoffman and Johnson, 2016; Tomczak and Welling, 2017) restricted on modalities from $\mathcal{S}$ and $q^{\star}(x_{\mathcal{S}}|z)=q_{\phi}(x_{\mathcal{S}},z)/q_{\phi}^{\text{agg}}(z)$ and likewise consider its conditional version,
@@
for an aggregated encoder conditioned on $x_{\mathcal{S}}$. We provide a multi-model ELBO surgery, summarized in Proposition 1 below. It implies that maximizing $\int p_{d}(x_{\mathcal{S}})\mathcal{L}_{\mathcal{S}}(x_{\mathcal{S}},\theta,\phi)\mathrm{d}x_{\mathcal{S}}$ drives
@@
$-\int p_{d}(x_{\setminus\mathcal{S}}|x_{\mathcal{S}})\left(\mathsf{K}\mathsf{L}(q_{\phi}(z|x)|p_{\theta}(z|x))+\int q_{\phi}(z|x)\log\frac{q_{\phi}(z|x_{\mathcal{S}})}{p_{\theta}(z|x_{\mathcal{S}})}\mathrm{d}z\right)\mathrm{d}x_{\setminus\mathcal{S}}$
@@
Our approach recovers meta-learning with (latent) Neural processes (Garnelo et al., 2018b) when one optimizes only $\mathcal{L}_{\setminus\mathcal{S}}$ with $\mathcal{S}$ determined by context-target splits, cf. Appendix B. Our analysis implies that $\mathcal{L}_{\mathcal{S}}+\mathcal{L}_{\setminus\mathcal{L}}$ is an approximate lower bound on the multi-modal log-likelihood that becomes tight for infinite-capacity encoders and is a true lower bound if $\mathsf{K}\mathsf{L}(q_{\phi,\setminus\mathcal{S}}^{\text{agg}}(z|x_{\mathcal{S}})|q_{\phi}(z|x_{\mathcal{S}}))=0$, see Remarks 3 and 5 for details.
Remark 3 (Log-Likelihood approximation, Empirical Bayes and Vamp-Priors)
@@
arising in Corollary 19 and in ($\mathsf{X}_{\text{conditional}}$) is not necessarily negative. Analogous to other variational approaches for learning conditional distributions such as latent Neural processes, our bound becomes an approximately lower bound. Note that $\mathcal{L}_{\mathcal{S}}$ is maximized when $q_{\phi}(z|x_{\mathcal{S}})=p_{\theta}(z|x_{\mathcal{S}})$, see ($\mathsf{X}_{\text{marginal}}$), while $\mathcal{L}_{\setminus\mathcal{S}}$ is maximized when
$q_{\phi}(z|x_{\mathcal{S}})=\int p_{d}(x_{\setminus\mathcal{S}}|x_{\mathcal{S}})q_{\phi}(z|x)\mathrm{d}x_{\setminus\mathcal{S}}=q_{\phi,\setminus\mathcal{S}}^{\text{agg}}(z|x_{\mathcal{S}}),$
see (1). The latter condition implies a lower bound in Corollary 19 of
@@
From a different perspective, we can consider an Empirical Bayes viewpoint (Robbins, 1992; Wang et al., 2019b) wherein one chooses the hyperparameters of the (conditional) prior so that it maximizes an approximation of the conditional log-likelihood $\log p_{\theta}(x_{\setminus\mathcal{S}}|x_{\mathcal{S}})$. The conditional prior $q_{\phi}(z|x_{\mathcal{S}})$ in the conditional ELBO term $\mathcal{L}_{\setminus\mathcal{S}}$ can thus be seen as a learned prior for any modality subset $\mathcal{S}$. While the aggregated prior $q_{\phi,\setminus\mathcal{S}}^{\text{agg}}(z|x_{\mathcal{S}})$ is the optimal learned prior when maximizing $\mathcal{L}_{\setminus\mathcal{S}}$, this choice can lead to overfitting. Following the approach of Tomczak and Welling (2017), one may consider learned priors in the form of a variational mixture of posteriors prior (VampPrior),
$p^{\text{Vamp}}(z|\mathcal{D})=\sum_{k=1}^{K}q_{\phi}(z|u_{k})$ (5)
for $K$ pseudo-inputs $u_{k}\in\mathsf{X}$ and which are learned end-to-end based on the full multi-modal empirical dataset $\mathcal{D}$ [^3] The aggregated posterior $q_{\phi,\setminus\mathcal{S}}^{\text{agg}}(z|x_{\mathcal{S}})$ also depends on the empirical data $p_{d}(x_{\setminus\mathcal{S}}|x_{\mathcal{S}})$ although this is not explicit in our notation. Indeed, Joy et al. (2021) considered a somewhat similar model for the conditional prior through mutual supervision. We contend ourselves with choosing the conditional prior as $q_{\phi}(z|x_{\mathcal{S}})$. Other choices are clearly possible, such as (5) leading to less regularised models, with algorithmic adjustment as in the uni-modal case. These choices also work with the flexible aggregation models introduced below. Likewise, one may consider VampPriors for the prior $p_{\theta}(z)$ in the bound $\mathcal{L}_{\mathcal{S}}$.
@@
In the uni-modal setting, the mismatch between the prior and the aggregated prior can be large and can lead to poor unconditional generative performance because this would lead to high-probability regions under the prior that have not been trained due to their small mass under the aggregated prior (Hoffman and Johnson, 2016; Rosca et al., 2018). Equation ($\mathsf{Z}_{\text{marginal}}$) extends this to the multi-modal case, and we expect that unconditional generation can be poor if this mismatch is large. Moreover, (1) extends this conditioned on some modality subset, and we expect that cross-generation for $x_{\setminus\mathcal{S}}$ conditional on $x_{\mathcal{S}}$ can be poor if the mismatch between $q_{\phi,\setminus\mathcal{S}}^{\text{agg}}(z|x_{\mathcal{S}})$ and $q_{\phi}(z|x_{\mathcal{S}})$ is large for $x_{\mathcal{S}}\sim p_{d}$, because high-probability regions under $q_{\phi}(z|x_{\mathcal{S}})$ will not have been trained - via optimizing $\mathcal{L}_{\setminus\mathcal{S}}(x)$ to model $x_{\setminus\mathcal{S}}$ conditional on $x_{\mathcal{S}}$, due to their small mass under $q_{\phi,\setminus\mathcal{S}}^{\text{agg}}(z|x_{\mathcal{S}})$. The mismatch will vanish when the encoders are consistent and correspond to a single Bayesian model where they approximate the true posterior distributions. A potential approach to reduce this mismatch may be to include as a regulariser the divergence between them that can be optimized by likelihood-free techniques, such as the Maximum-Mean Discrepancy (Gretton et al., 2006), as in Zhao et al. (2019) for uni-modal or unconditional models.
@@
Corollary 19 shows that the variational bound can become tight in the limiting case where the encoding distributions approximate the true posterior distributions. A similar result does not hold for the mixture-based multi-modal bound. Indeed, as shown in Daunhawer et al. (2022), there is a gap between the variational bound and the log-likelihood given by the conditional entropies that cannot be reduced even for flexible encoding distributions. More precisely, it holds that
$\int p_{d}(x)\log p_{\theta}(x)\mathrm{d}x\geq\int p_{d}(x)\mathcal{L}^{\text{Mix}}(x,\theta,\phi,1)\mathrm{d}x+\mathcal{H}(p_{d}(X_{\setminus\mathcal{S}}|X_{\mathcal{S}})).$
Moreover, our bound can be tight for an arbitrary number of modalities in the limiting case of infinite-capacity encoders. In contrast, Daunhawer et al. (2022) show that for mixture-based bounds, this variational gap increases with each additional modality if the new modality is ’sufficiently diverse’, even for infinite-capacity encoders. Note that in practice, we optimize over the objective where the mask $\mathcal{S}$ is not fixed but random, which induces a Jensen gap as in other any-order methods (Hoogeboom et al., 2021; Shih et al., 2022).
@@
For simplicity, we have chosen to sample $\mathcal{S}\sim\rho$ in our experiments via the hierarchical construction $\gamma\sim\mathcal{U}(0,1)$, $m_{j}\sim\text{Bern}(\gamma)$ iid for all $j\in[M]$ and setting $\mathcal{S}=\{s\in[M]\colon m_{j}=1\}$. The distribution $\rho$ for masking the modalities can be adjusted to accommodate various weights for different modality subsets. Indeed, (2) can be seen as a linear scalarization of a multi-task learning problem (Fliege and Svaiter, 2000; Sener and Koltun, 2018). We aim to optimize a loss vector $(\mathcal{L}_{\mathcal{S}}+\mathcal{L}_{\setminus\mathcal{S}})_{\mathcal{S}\subset\mathcal{M}}$, where the gradients for each $\mathcal{S}\subset\mathcal{M}$ can point in different directions, making it challenging to minimize the loss for all modalities simultaneously. Consequently, Javaloy et al. (2022) used multi-task learning techniques (e.g., as suggested in Chen et al. (2018); Yu et al. (2020)) for adjusting the gradients in mixture-based VAEs. Such improved optimization routines are orthogonal to our approach. Similarly, we do not analyze optimization issues such as initializations and training dynamics that have been found challenging for multi-modal learning (Wang et al., 2020; Huang et al., 2022).

### v3

@@ 2 A tighter variational objective with arbitrary modality masking
This is simply a standard variational lower bound [Jordan et al., 1999, Blei et al., 2017] restricted to the subset $\mathcal{S}$ for $\beta=1$, and therefore $\mathcal{L}_{\mathcal{S}}(x_{\mathcal{S}},\theta,\phi,1)\leq\log p_{\theta}(x_{\mathcal{S}})$. One can express the variational bound in information-theoretic [Alemi et al., 2018] terms as
$\int p_{d}(x_{\mathcal{S}})\mathcal{L}(x_{\mathcal{S}})\mathrm{d}x_{\mathcal{S}}=-D_{\mathcal{S}}-\beta R_{\mathcal{S}}$
for the rate
measuring the information content that is encoded from the observed modalities in $\mathcal{S}$ by $q_{\phi}$ into the latent representation, and the distortion
$D_{\mathcal{S}}=-\int p_{d}(x_{\mathcal{S}})q_{\phi}(z|x_{\mathcal{S}})\log p_{\theta}(x_{\mathcal{S}}|z)\mathrm{d}z\mathrm{d}x_{\mathcal{S}}$
given as the negative reconstruction log-likelihood of the modalities in $\mathcal{S}$. While the latent variable $Z_{\mathcal{S}}$ that is encoded via $q_{\phi}(z|x_{\mathcal{S}})$ from $X_{\mathcal{S}}$ can be tuned via the choice of $\beta>0$ to tradeoff compression and reconstruction of all modalities in $\mathcal{S}$ jointly, it does not explicitly optimize for cross-modal prediction of modalities not in $\mathcal{S}$. Indeed, the mixture-based variational bound differs from the above decomposition exactly by an additional cross-modal prediction or cross-distortion term
$D^{c}_{\setminus\mathcal{S}}=-\int p_{d}(x)q_{\phi}(z_{\mathcal{S}}|x_{\mathcal{S}})\log p_{\theta}(x_{\setminus\mathcal{S}}|z_{\mathcal{S}})\mathrm{d}z_{\mathcal{S}}\mathrm{d}x,$
thereby explicitly optimizing for both self-reconstruction of a modality subset and cross-modal prediction within a single objective:
$\int p_{d}(\mathrm{d}x)\mathcal{L}^{\text{Mix}}_{\mathcal{S}}(x)=-D_{\mathcal{S}}-D^{c}_{\setminus\mathcal{S}}-\beta R_{\mathcal{S}}.$ (4)
Instead of adding an explicit cross-modal prediction term, we consider an additional variational objective with a second latent variable $Z_{\mathcal{M}}$ that encodes all observed modalities $X=X_{\mathcal{M}}$ and tries to reconstruct the remaining modality subset. Unlike the latent variable $Z_{\mathcal{S}}$ in (1) and (3) that can only encode incomplete information using the modalities $\mathcal{S}$, this second latent variable $Z_{\mathcal{M}}$ can encode modality-specific information from all observed modalities, thereby avoiding the averaging prediction in the mixture-based bound, see Figure 1(b).
Ideally, we may like to consider an additional variational objective that lower bounds the conditional log-likelihood $\log p_{\theta}(x_{\setminus\mathcal{S}}|x_{\mathcal{S}})$ so that maximizing the sum of both bounds maximizes a lower bound of the multi-modal log-likelihood $\log p_{\theta}(x_{\mathcal{S}},x_{\setminus\mathcal{S}})=\log p_{\theta}(x_{\mathcal{S}})+\log p_{\theta}(x_{\setminus\mathcal{S}}|x_{\mathcal{S}})$. To motivate such a conditional variational objective, note that maximizing the variational bound (3) with infinite capacity encoders yields $q_{\phi}(z|x_{\mathcal{S}})=p_{\theta}(z|x_{\mathcal{S}})$. This suggests replacing the intractable posterior $p_{\theta}(z|x_{\mathcal{S}})$ with the encoder $q_{\phi}(z|x_{\mathcal{S}})$ for the probabilistic model when conditioned on $x_{\mathcal{S}}$. A variational objective under this replacement then becomes
@@
However, the above only approximates a lower bound on the conditional log-likelihood $\log p_{\theta}(x_{\setminus\mathcal{S}}|x_{\mathcal{S}})$, and $\mathcal{L}_{\setminus\mathcal{S}}$ is a lower bound only under idealized conditions. We will make these approximations more precise in Section 2.1, where we illustrate how these bounds yield to a matching of different distributions in the latent or data space, while Section 2.2 provides an information-theoretic interpretation of these variational objectives. The introduction of a second latent variable leads to two KL regularization terms, see Figure 1(d), that do not satisfy a triangle inequality $\mathsf{K}\mathsf{L}(q_{\phi}(z|x)|q_{\phi}(z|x_{\mathcal{S}}))+\mathsf{K}\mathsf{L}(q_{\phi}(z|x_{\mathcal{S}})|p_{\theta}(z))\ngeq\mathsf{K}\mathsf{L}(q_{\phi}(z|x)|p_{\theta}(z))$.
In summary, for some fixed density $\rho$ on $\mathcal{P}(\mathcal{M})$, we suggest to maximize the overall bound
@@
with respect to $\theta$ and $\phi$, which is a generalization of the bound suggested in Wu and Goodman [2019] to an arbitrary number of modalities. This bound can be optimized using standard Monte Carlo techniques, for example, by computing unbiased pathwise gradients [Kingma and Ba, 2014, Rezende et al., 2014, Titsias and Lázaro-Gredilla, 2014] using the reparameterization trick. For variational families such as Gaussian mixtures [^2] For MoE aggregation schemes, Shi et al. [2019] considered a stratified ELBO estimator as well as a tighter bound based on importance sampling, see also Morningstar et al. [2021], that we do not pursue here for consistency with other aggregation schemes that can likewise be optimized based on importance sampling ideas., one can employ implicit reparameterization [Figurnov et al., 2018]. It is straightforward to adapt variance reduction techniques such as ignoring the scoring term of the multi-modal encoding densities for pathwise gradients [Roeder et al., 2017], see Algorithm 1 in Appendix H for pseudo-code. Nevertheless, a scalable approach requires an encoding technique that allows to condition on any masked modalities with a computational complexity that does not increase exponentially in $M$. We will analyse scalable model architectures in Section 3.
@@ 2.1 Multi-modal distribution matching
Likelihood-based learning approaches aim to match the model distribution $p_{\theta}(x)$ to the true data distribution $p_{d}(x)$. Variational approaches achieve this by matching in the latent space the encoding distribution to the true posterior as well as maximizing a tight lower bound on $\log p_{\theta}(x)$, see Rosca et al. [2018]. These types of analyses have proved useful for uni-modal VAEs as they can provide some insights as to why VAEs may lead to worse generative sample quality compared to other generative models such as GANs [Goodfellow et al., 2014] or may fail to learn useful latent representations [Zhao et al., 2019, Dieng et al., 2019]. We show similar results for the multi-modal variational objectives. This suggests that limitations from uni-modal VAEs also affect multi-modal VAEs, but also that previous attempts to address these shortcomings in uni-modal VAEs may benefit multi-modal VAEs. In particular, mismatches between the prior and the aggregated prior for uni-modal VAEs that result in poor unconditional generation have a natural counterpart for cross-modal generations with multi-modal VAEs that may potentially be reduced using more flexible conditional prior distributions, see Remark 3, or via adding additional mutual information regularising terms [Zhao et al., 2019], see Remark 4. Given these results, it is neither surprising that multi-modal diffusion models such as Bounoua et al. [2023], Bao et al. [2023] yield improved sample quality, nor that sample quality can be improved by augmenting multi-modal VAEs with diffusion models [Pandey et al., 2022, Palumbo et al., 2024].
We consider the densities
@@
for the aggregated prior [Makhzani et al., 2016, Hoffman and Johnson, 2016, Tomczak and Welling, 2017] restricted on modalities from $\mathcal{S}$ and $q^{\star}(x_{\mathcal{S}}|z)=q_{\phi}(x_{\mathcal{S}},z)/q_{\phi}^{\text{agg}}(z)$ and likewise consider its conditional version,
@@
for an aggregated encoder conditioned on $x_{\mathcal{S}}$. We provide a multi-modal ELBO surgery, summarized in Proposition 1 below. It implies that maximizing $\int p_{d}(x_{\mathcal{S}})\mathcal{L}_{\mathcal{S}}(x_{\mathcal{S}},\theta,\phi)\mathrm{d}x_{\mathcal{S}}$ drives
@@
$-\int p_{d}(x_{\setminus\mathcal{S}}|x_{\mathcal{S}})\left(\mathsf{K}\mathsf{L}(q_{\phi}(z|x)|p_{\theta}(z|x))-\int q_{\phi}(z|x)\log\frac{q_{\phi}(z|x_{\mathcal{S}})}{p_{\theta}(z|x_{\mathcal{S}})}\mathrm{d}z\right)\mathrm{d}x_{\setminus\mathcal{S}}$
@@
Our approach recovers meta-learning with (latent) Neural processes [Garnelo et al., 2018b] when one optimizes only $\mathcal{L}_{\setminus\mathcal{S}}$ with $\mathcal{S}$ determined by context-target splits, cf. Appendix B. Our analysis implies that $\mathcal{L}_{\mathcal{S}}+\mathcal{L}_{\setminus\mathcal{S}}$ is an approximation of a lower bound on the multi-modal log-likelihood that becomes tight for infinite-capacity encoders so that $q_{\phi}(z|x_{\mathcal{S}})=p_{\theta}(z|x_{\mathcal{S}})$ and $q_{\phi}(z|x)=p_{\theta}(z|x)$, see Remarks 3 and 5 for details.
Remark 3 (Log-Likelihood approximation and Empirical Bayes)
@@
arising in Corollary 2 and in ($\mathsf{X}_{\text{conditional}}$) is not necessarily negative. Analogous to other variational approaches for learning conditional distributions such as latent Neural processes, our bound becomes an approximation of a lower bound. Note that $\mathcal{L}_{\mathcal{S}}$ is maximized when $q_{\phi}(z|x_{\mathcal{S}})=p_{\theta}(z|x_{\mathcal{S}})$, see ($\mathsf{X}_{\text{marginal}}$), which implies a lower bound in Corollary 2 of
@@
We can re-write the conditional expectation of $\mathcal{L}_{\setminus\mathcal{S}}$ for any fixed $x_{\mathcal{S}}$ as
$\int p_{d}(x_{\setminus\mathcal{S}}|x_{\mathcal{S}})\mathcal{L}_{\setminus\mathcal{S}}(x,\theta,\phi,1)\mathrm{d}x_{\setminus\mathcal{S}}=$ $\int p_{d}(x_{\setminus\mathcal{S}}|x_{\mathcal{S}})q_{\phi}(z|x)\log p_{\theta}(x_{\setminus\mathcal{S}}|z)\mathrm{d}z\mathrm{d}x_{\setminus\mathcal{S}}+p_{d}(x_{\setminus\mathcal{S}}|x_{\mathcal{S}})\mathcal{H}(q_{\phi}(z|x))\mathrm{d}x_{\setminus\mathcal{S}}$
$+\int q_{\phi,\setminus\mathcal{S}}^{\text{agg}}(z|x_{\mathcal{S}})\log q_{\phi}(z|x_{\mathcal{S}})\mathrm{d}z.$
Whenever $q_{\phi}(z|x_{\mathcal{S}})$ can be learned independently from $q_{\phi}(z|x)$, the above is maximized for
$q_{\phi}(z|x_{\mathcal{S}})=\int p_{d}(x_{\setminus\mathcal{S}}|x_{\mathcal{S}})q_{\phi}(z|x)\mathrm{d}x_{\setminus\mathcal{S}}=q_{\phi,\setminus\mathcal{S}}^{\text{agg}}(z|x_{\mathcal{S}}).$
From a different perspective, we can consider an Empirical Bayes viewpoint [Robbins, 1992, Wang et al., 2019b] wherein one chooses the hyperparameters of the (conditional) prior so that it maximizes an approximation of the conditional log-likelihood $\log p_{\theta}(x_{\setminus\mathcal{S}}|x_{\mathcal{S}})$. The conditional prior $p^{\star}_{\vartheta}(z|x_{\mathcal{S}})$ in the corresponding conditional ELBO term
$\mathcal{L}^{\star}_{\setminus\mathcal{S}}(x,\theta,\phi,\vartheta,\beta)=\int q_{\phi}(z|x)\left[\log p_{\theta}(x_{\setminus\mathcal{S}}|z)\right]\mathrm{d}z-\beta\mathsf{K}\mathsf{L}(q_{\phi}(z|x)|p^{\star}_{\vartheta}(z|x_{\mathcal{S}})).$ (6)
can thus be seen as a learned prior having the parameter $\vartheta=\vartheta(\mathcal{D})$ that is learned by maximizing the above variational approximation of $\log p_{\theta}(x_{\setminus\mathcal{S}}|x_{\mathcal{S}})$ over $x\sim p_{d}$ for $\beta=1$, and as such depends on the empirical multi-modal dataset $\mathcal{D}$. While the aggregated prior $q_{\phi,\setminus\mathcal{S}}^{\text{agg}}(z|x_{\mathcal{S}})$ is the optimal learned prior when maximizing $\mathcal{L}_{\setminus\mathcal{S}}$, this choice can lead to overfitting. Moreover, computation of the aggregated prior, or sparse approximations thereof, such as variational mixture of posteriors prior (VampPrior) in Tomczak and Welling [2017], are challenging in the conditional setup. Previous constructions in this direction [Joy et al., 2021] for learning priors for bi-modal data only considered unconditional versions, wherein pseudo-samples are not dependent on some condition $x_{\mathcal{S}}$. While our permutation-invariant architectures introduced below may be used for flexibly parameterizing conditional prior distributions $p^{\star}_{\vartheta}(z|x_{\mathcal{S}})$ as a function of $x_{\mathcal{S}}$ with a model that is different from the encoding distributions $q_{\phi}(z|x_{\mathcal{S}})$, we contend ourselves with choosing the same model for both the conditional prior and the encoding distribution, $p^{\star}_{\vartheta}(z|x_{\mathcal{S}})=q_{\phi}(z|x_{\mathcal{S}})$. Note that the encoding distribution then features both bounds, which encourages learning encoding distributions that perform well as conditional priors and as encoding distributions. In the ideal scenario where both the generative and inference models have the flexibility to satisfy $p_{d}(x_{\setminus\mathcal{S}}|x_{\mathcal{S}})=p_{\theta}(x_{\setminus\mathcal{S}}|x_{\mathcal{S}})$, $q_{\phi}(z|x_{\mathcal{S}})=p_{\theta}(z|x_{\mathcal{S}})$ and $q_{\phi}(z|x_{\mathcal{S}})=p_{\theta}(z|x_{\mathcal{S}})$, then the optimal conditional prior distribution is
$q_{\phi,\setminus\mathcal{S}}^{\text{agg}}(z|x_{\mathcal{S}})=\int p_{\theta}(x_{\setminus\mathcal{S}}|x_{\mathcal{S}})p_{\theta}(z|x)\mathrm{d}x_{\setminus\mathcal{S}}=\int p_{\theta}(x_{\setminus\mathcal{S}}|x_{\mathcal{S}})\frac{p_{\theta}(z|x_{\mathcal{S}})p_{\theta}(x_{\setminus\mathcal{S}}|x_{\mathcal{S}})}{p_{\theta}(x_{\setminus\mathcal{S}}|x_{\mathcal{S}})}\mathrm{d}x_{\setminus\mathcal{S}}=p_{\theta}(z|x_{\mathcal{S}})=q_{\phi}(z|x_{\mathcal{S}}).$
@@
In the uni-modal setting, the mismatch between the prior and the aggregated prior can be large and can lead to poor unconditional generative performance because this would lead to high-probability regions under the prior that have not been trained due to their small mass under the aggregated prior [Hoffman and Johnson, 2016, Rosca et al., 2018]. Equation ($\mathsf{Z}_{\text{marginal}}$) extends this to the multi-modal case, and we expect that unconditional generation can be poor if this mismatch is large. Moreover, (1) extends this conditioned on some modality subset, and we expect that cross-generation for $x_{\setminus\mathcal{S}}$ conditional on $x_{\mathcal{S}}$ can be poor if the mismatch between $q_{\phi,\setminus\mathcal{S}}^{\text{agg}}(z|x_{\mathcal{S}})$ and $q_{\phi}(z|x_{\mathcal{S}})$ is large for $x_{\mathcal{S}}\sim p_{d}$, because high-probability regions under $q_{\phi}(z|x_{\mathcal{S}})$ will not have been trained - via optimizing $\mathcal{L}_{\setminus\mathcal{S}}(x)$ to model $x_{\setminus\mathcal{S}}$ conditional on $x_{\mathcal{S}}$, due to their small mass under $q_{\phi,\setminus\mathcal{S}}^{\text{agg}}(z|x_{\mathcal{S}})$. The mismatch will vanish when the encoders are consistent and correspond to a single Bayesian model where they approximate the true posterior distributions. A potential approach to reduce this mismatch may be to include as a regulariser the divergence between them that can be optimized by likelihood-free techniques, such as the Maximum-Mean Discrepancy [Gretton et al., 2006], as in Zhao et al. [2019] for uni-modal or unconditional models. For the mixture-based bound, the same distribution mismatch affects unconditional generation, while both the training and generative sampling distribution is $q_{\phi}(z|x_{\mathcal{S}})$ for cross-generation.
@@
Corollary 2 shows that the variational objective can become a tight bound in the limiting case where the encoding distributions approximate the true posterior distributions. A similar result does not hold for the mixture-based multi-modal bound. Moreover, our bound can be tight for an arbitrary number of modalities in the limiting case of infinite-capacity encoders. In contrast, Daunhawer et al. [2022] show that for mixture-based bounds, this variational gap increases with each additional modality if the new modality is ’sufficiently diverse’, even for infinite-capacity encoders.
@@
For simplicity, we have chosen to sample $\mathcal{S}\sim\rho$ in our experiments via the hierarchical construction $\gamma\sim\mathcal{U}(0,1)$, $m_{j}\sim\text{Bern}(\gamma)$ iid for all $j\in[M]$ and setting $\mathcal{S}=\{s\in[M]\colon m_{j}=1\}$. The distribution $\rho$ for masking the modalities can be adjusted to accommodate various weights for different modality subsets. Indeed, (2) can be seen as a linear scalarization of a multi-task learning problem [Fliege and Svaiter, 2000, Sener and Koltun, 2018]. We aim to optimize a loss vector $(\mathcal{L}_{\mathcal{S}}+\mathcal{L}_{\setminus\mathcal{S}})_{\mathcal{S}\subset\mathcal{M}}$, where the gradients for each $\mathcal{S}\subset\mathcal{M}$ can point in different directions, making it challenging to minimize the loss for all modalities simultaneously. Consequently, Javaloy et al. [2022] used multi-task learning techniques (e.g., as suggested in Chen et al. [2018], Yu et al. [2020]) for adjusting the gradients in mixture-based VAEs. Such improved optimization routines are orthogonal to our approach. Similarly, we do not analyze optimization issues such as initializations and training dynamics that have been found challenging for multi-modal learning [Wang et al., 2020, Huang et al., 2022].